For the hydrolysis of a salt of a weak acid and a weak base, show that
Kh=Kw/KaKb
1
Expert's answer
2020-04-15T01:52:22-0400

salt hydrolysis of weak acid and weak base

"CH_3COOH+NH_4OH\to CH_3COO^-NH_4^++H_2O"

hydrolysis of salt

"CH_3COO^-NH_4^++H_2O\to CH_3COOH+NH_4OH"

"K_h=\frac{CH_3COOH][NH_4OH]}{[CH_3COO^-NH_4^+]}"

"CH_3COOH\to CH_3COO^-+H^+"

"K_a=\frac{[CH_3COO^-][H^+]}{[CH_3COOH]}"

"NH_4OH\to NH_4^++OH^-"

"K_b=\frac{[NH_4^+][OH^-]}{[NH_4OH]}"

"H_2O\to H^++OH^-"

"K_w=[H^+][OH^-]"

Hence,

"K_h=\frac{K_w}{K_aK_b}"